# TimeSieve: Extracting Temporal Dynamics through Information Bottleneck

### v1

@@ Abstract.
Time series forecasting has become an increasingly popular research area due to its critical applications in various real-world domains such as traffic management, weather prediction, and financial analysis. Despite significant advancements, existing models face notable challenges, including the necessity of manual hyperparameter tuning for different datasets, and difficulty in effectively distinguishing signal from redundant features in data characterized by strong seasonality. These issues hinder the generalization and practical application of time series forecasting models. To solve this issues, we propose an innovative time series forecasting model - TimeSieve designed to address these challenges. Our approach employs wavelet transforms to preprocess time series data, effectively capturing multi-scale features without the need for additional parameters or manual hyperparameter tuning. Additionally, we introduce the information bottleneck theory that filters out redundant features from both detail and approximation coefficients, retaining only the most predictive information. This combination reduces significantly improves the model’s accuracy. Extensive experiments demonstrate that our model outperforms existing state-of-the-art methods on 70% of the datasets, achieving higher predictive accuracy and better generalization across diverse datasets. Our results validate the effectiveness of our approach in addressing the key challenges in time series forecasting, paving the way for more reliable and efficient predictive models in practical applications. The code for our model is available at https://github.com/xll0328/TimeSieve.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
[…]
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
@@
The field of time series forecasting has experienced significant advancements and found widespread applications in various real-world engineering tasks, such as traffic management (Lana et al., 2018), pandemic spread (Kumar and Susan, 2020), and financial analysis (Sezer et al., 2020). From early models like GRU (Dey and Salem, 2017) and LSTM (Greff et al., 2016) to the more recent Transformer-based architectures (Liu et al., 2023b; Wu et al., 2021), a plethora of sophisticated models have emerged. For instance, TimesNet (Wu et al., 2023) and FreTS (Yi et al., 2023) approach data processing from a frequency domain perspective, while multi-scale models like Pyraformer (Liu et al., 2022) and Scaleformer (Shabani et al., 2023) leverage intricate multi-scale and cross-dimensional dependency structures to significantly enhance performance on complex forecasting tasks.
@@
Howerver, existing models often require the introduction of additional parameters to capture multi-scale features, which increases computational complexity. Moreover, these models necessitate manual tuning of hyperparameters for different datasets, further complicating their application and limiting their generalization across diverse scenarios. More importantly, many models, including Autoformer (Wu et al., 2021) , struggle to effectively distinguish between signal and redundant features when dealing with data characterized by strong seasonality. Redundant features not only interferes with the learning process, leading to increased errors, but also obscures the true patterns within the data, making the predictions unreliable. This issue is particularly pronounced when traditional autocorrelation mechanisms break down due to redundant features interference, significantly diminishing the practical value of these models.
@@
To address these challenges, we proposes an innovative time series forecasting model designed to significantly improve predictive accuracy and broaden applicability. First, we employ wavelet transforms to preprocess the time series data using the Wavelet Decomposition Block (WDB) and the Wavelet Reconstruction Block (WRB). Wavelet transforms efficiently capture information from multiple scales, but this process may also introduce redundant features. To address this, we integrate the Information Filtering and Compression Block (IFCB), an information bottleneck module that compresses and extracts the most relevant features by filtering out redundant features from both detail and approximation coefficients. These components work together, significantly enhancing the model’s accuracy.
@@
  • We introduce WDB and WRB that efficiently capture multi-scale features in time series data without requiring manual hyperparameter tuning or the introduction of additional parameters.
@@
  • We utilize an information-theoretic approach by incorporating an IFCB to filter out redundant features from the data. This ensures that only the most predictive features are retained, enhancing the model’s accuracy.
@@
  • By integrating wavelet transforms for comprehensive feature extraction and the information bottleneck block for redundant features, our model achieves state-of-the-art performance on 70% of the datasets tested, demonstrating superior predictive accuracy and generalization across diverse scenarios.
The rest of this paper is organized as follows. Section 2 reviews related work in the field of time series forecasting. In Section 3, we provide a detailed description of our model implementation. The experimental results and analysis are discussed in Section 4. Finally, Section 5 concludes the paper with a summary of our findings and suggestions for future research.
2. Related Work
2.1. Time Series Forecasting
@@
In recent years, the field of time series forecasting has advanced rapidly, giving rise to many efficient models. These models predominantly rely on multilayer perceptron (MLP) approaches like MSD-Mixer (Zhong et al., 2024), TSMixer (Chen et al., 2023),DLinear (Zeng et al., 2022)and FreTs (Yi et al., 2023), along with Transformer-based techniques such as PatchTST (Nie et al., 2023), Autoformer, and Informer (Zhou et al., 2021), to perform predictions using sophisticated data transformation and learning mechanisms.
Time series forecasting can be defined as follows: Given a historical sequence of observations $\ X=\{x_{1},x_{2},\ldots,x_{T}\}$, where $T$ is the lookback window length, the goal is to predict future values $\ Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}$, where $H$ is the forecast window length.
(1) $Y=F(X)$
@@
where $F$ is the predictive model mapping the input sequence $X$ to the output sequence $Y$. However, time series data often exhibit complex, multi-scale characteristics. Conventional feature extraction methods, like standard pooling, require additional parameter adjustments or rely on manually designed pooling sizes, increasing model complexity and limiting flexibility.
Wavelet transform, a classical signal processing technique (Farge, 1992), excels in multi-scale analysis. It decomposes signals into different frequency components, aiding in the understanding and prediction of intricate patterns.
@@
We proposes a novel approach that combines wavelet transform with modern machine learning techniques for time series forecasting. The wavelet transform preprocesses the input data, decomposing it into detail coefficients and approximation coefficients at different scales. We believe this method effectively captures comprehensive features from the time series data.
2.2. Information Bottleneck Theory
@@
In recent years, with the growing intersection of information theory and deep learning research, many researchers have begun analyzing and optimizing neural networks from the perspective of information theory. In this trend, Information Bottleneck (IB) theory has gained particular attention (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017), finding wide application in emerging fields such as graph neural networks (Dai et al., 2023), multimodal learning (Mai et al., 2022), and recently in time series imputation (Choi and Lee, 2024). IB theory helps improve model performance by optimizing information flow and processing (Tishby and Zaslavsky, 2015).
@@
We believe that wavelet transformation can effectively decompose time series data into components that contain valuable information. However, these components also include a significant amount of redundant information. If not properly addressed, this redundant information can interfere with the model’s learning performance and predictive accuracy. From the perspective of information theory, IB theory provides a mechanism to effectively retain critical information related to target variables while filtering out irrelevant or redundant information in the model’s intermediate representations. This approach enhances the model’s ability to capture key features. Furthermore, it strikes a delicate balance between compressing input information and maintaining predictive accuracy, ensuring that the model’s output remains precise and concise by:
(2) $\max I(y;z)$
@@
3. TimeSieve
Figure 1. Overview of the TimeSieve Framework. The upper part illustrates the overall architecture of TimeSieve, which consists of the Wavelet Decomposition Block (WDB), Wavelet Reconstruction Block (WRB), Information Fusion and Compression Block (IFCB), and the prediction. The input time series data, $T\times C$, is decomposed by WDB into coefficients $cA$ and $cD$, which are processed by IFCB. The WRB then reconstructs the data, followed by the predictor generating the corresponding forecast steps. The lower part provides a detailed visualization of the data flow within WDB and IFCB, with further visualization of IFCB available in Figure 2.
In our study, we propose a novel time series forecasting model named TimeSieve. This model integrates the Information Filtering and Compression Block (IFCB) with wavelet transform technology to enhance the accuracy of time series predictions.
@@ 3. TimeSieve
(3) $[cA,cD]=WDB(X)$
where the approximation coefficients $cA$ representing the low-frequency trends, and the detail coefficients $cD$ representing the high-frequency details.
However, the extracted multi-scale information may contain redundant features, which can adversely affect the model’s learning performance and predictive accuracy. To address this issue, we introduce the Information Filtering and Compression Block (IFCB) to filter the information. Additionally, we employ residual connections to ensure that valuable information is not lost during the filtering process. The IFCB optimizes the information flow, retaining critical information while filtering out irrelevant or redundant information. The equations for the Information Bottleneck module are as follows:
(4) $c\hat{A}=IFCB(cA)+cA$
(5) $c\hat{D}=IFCB(cD)+cD$
where $c\hat{A}$ and $c\hat{D}$ represents the filtered coefficients $cA$ and $cD$. After filtering with the IFCB, we apply the Wavelet Reconstruction Block (WRB) to reconstruct the processed data back into the time domain. This step ensures that the features at different scales are fully utilized. Finally, we use a Multi-Layer Perceptron (MLP) as the prediction to make the final forecast on the processed time series data. The equation for the prediction step is as follows:
(6) $Y=MLP(WRB([c\hat{A},c\hat{D}]))$
@@
3.1. WDB and WRB
To effectively capture comprehensive feature information, we employ wavelet transformation to decompose time series data inspired by (Meyer et al., 2023; Nason and Wei, 2024), allowing us to analyze the data across multiple scales. Wavelet transformation allows us to break down the data into various frequency components, enabling us to analyze and extract features across multiple temporal scales. This multi-scale analysis is crucial for identifying patterns and trends that occur at different time resolutions.
WDB. For the input time series $x(t)\in\mathbb{R}^{T\times C}$, where $T$ represents the length of the lookback window and $C$ denotes the number of variates, the wavelet transform decomposes the series into approximation coefficients $cA$ and detail coefficients $cD$. The approximation coefficients $cA$ capture the general trend or smooth component of the series, representing low-frequency information. In contrast, the detail coefficients $cD$ capture the fine-grained, high-frequency information, highlighting local fluctuations and details within the series. This decomposition can be expressed as:
[…]
@@ 3.1. WDB and WRB
(9) $\phi(x)=\sum_{k=0}^{N-1}a_{k}\phi(2u-k),$
(10) $\psi(x)=\sum_{k=0}^{M-1}b_{k}\phi(2u-k),$
In the case of db1 wavelet transform, $\phi(t)$ represents the low-pass filter, capturing the smooth, approximate part of the signal, and $\psi(t)$ represents the high-pass filter, capturing the detailed, high-frequency part of the signal. Here, $a_{k}$ and $b_{k}$ are the filter coefficients for the scaling and wavelet functions, respectively, $u$ is the variable over which the functions are defined, and $k$ is the index of the coefficients.
@@
where $c\hat{A}$ and $c\hat{D}$ are the filtered approximation and detail coefficients, respectively. Finally, the reconstructed time series is passed through a MLP for final prediction.
3.2. IFCB
Motivated by the need to effectively filter out redundant features while retaining critical information, we introduce the Information Filtering and Compression Block (IFCB), inspired by the information bottleneck principle (Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017; Liu et al., 2024). In this section, we process the detail coefficients $cD$ and approximation coefficients $cA$ from the wavelet transform of the time series data, for convenience in subsequent explanations, we define $cI\in(cA,cD)$ to collectively represent $cA$ and $cD$. Additionally, We use $c\hat{I}$, where $I$ is in the range $[A,D]$. to represent $c\hat{A}$ and $c\hat{D}$, the filtered versions of $cA$ and $cD$. Initially, we define the input to the IFCB as $cI\in\mathbb{R}^{T/2\times C}$ and the output to the IFCB as $c\hat{I}\in\mathbb{R}^{T/2\times C}$.
In our model, both $cI$ and $c\hat{I}$ are treated as random variables with a joint distribution $p(i,\hat{i})$. The mutual information between them, denoted as $I(cI;c\hat{I})$, quantifies the amount of information shared by $cI$ and $c\hat{I}$ (More details you can see in Appendix B). It is defined as:
(12) $1(cI;c\hat{I})$ $=D_{kL}[p(i,\hat{i})\|p(i)p(\hat{i})]$
We define the intermediate hidden layer random variable $Z$. Following the principles of a Markov chain where $cI\rightarrow Z\rightarrow c\hat{I}$, the mutual information satisfies $I(cI;Z)\geq I(cI;c\hat{I})$ (More details see in Appendix A). Although there is an inevitable information loss in the transmission to the middle layer $Z$, the key is that $Z$ retains the most critical information for predicting $c\hat{I}$ while filtering out irrelevant features.
This allows us to define the objective of IFCB, which is to maximize the mutual information between $Z$ and $c\hat{I}$ while minimizing the information between $Z$ and $cI$.
(13) $\min\{I(cI;Z)-\beta I(Z;c\hat{I})\}$
where $\beta$ is a trade-off parameter that determines the weight of the minimal information constraint in the optimization process.
The optimization is carried out using a deep neural network, which is designed to minimize the above objective by adjusting the network weights and biases through backpropagation and suitable regularization techniques.
Figure 2. Overview of IFCB: This module consists of a reparameterization step and deep neural networks. Initially, the standard deviation and mean ($\mu_{z},\Sigma_{z}$) of the data are computed. The reparameterization is then performed using the equation $z=\mu_{z}+\Sigma_{z}\times\epsilon$. The input coefficients $cI$ are processed through the first neural network to obtain $\mu_{z}$ and $\Sigma_{z}$. The reparameterized $z$ is then fed into the second neural network to produce the reconstructed coefficients $\hat{cI}$.
Optimization is then performed using a deep neural network. Assuming $p(z|i)$ follows a Gaussian distribution, we can derive the following:
(14) $p(z|i)$ $=\mathcal{N}(\mu(i;\theta_{\mu}),\Sigma(i;\theta_{\Sigma}))$
$=\mathcal{N}(\mu_{z},\Sigma_{z})$
Parameter updates pose a challenge due to the stochastic nature of the gradient calculations. To address this, the reparameterization trick is introduced. The formula is given as:
(15) $z=\mu_{z}+\Sigma_{z}\times\epsilon$
where $\epsilon$ is sampled from a standard Gaussian distribution.
Given that the study focuses on a regression task, we define the conditional probability $q(\hat{i}|z)$ as follows:
(16) $q(\hat{i}|z)=e^{-||\hat{i}-D(z;\theta_{c})||}+C$
where $D$ is a decoder function parameterized by $\theta_{c}$, and $C$ is a constant. This formulation measures the fit of the model’s predictions by capturing the deviation of the predicted output $\hat{i}$ from the output of the decoder.
@@ 3.2. IFCB
We now specify the model’s loss function. This function is crucial for training the model to minimize prediction errors and optimize the distribution parameters effectively. The loss is composed of the original loss component and the IFCB loss, as shown below:
(17) $\mathcal{L}$ $=\mathcal{L}_{o}+\mathcal{L}_{IB}$
$=\mathcal{L}_{o}+D_{KL}[\mathcal{N}(\mu_{z},\Sigma_{z})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z)\,||\,p(z|i)]$

### v2

# TimeSieve: Extracting Temporal Dynamics via Information Bottleneck

@@ Abstract
Time series forecasting has become an increasingly popular research area due to its critical applications in various real-world domains such as traffic management, weather prediction, and financial analysis. Despite significant advancements, existing models face notable challenges, including the necessity of manual hyperparameter tuning for different datasets, and difficulty in effectively distinguishing signal from redundant features in data characterized by strong seasonality. These issues hinder the generalization and practical application of time series forecasting models. To solve this issues, this paper propose an innovative time series forecasting model - TimeSieve designed to address these challenges. Our approach employs wavelet transforms to preprocess time series data, effectively capturing multi-scale features without the need for additional parameters or manual hyperparameter tuning. Additionally, this paper introduce the information bottleneck theory that filters out redundant features from both detail and approximation coefficients, retaining only the most predictive information. This combination reduces significantly improves the model’s accuracy. Extensive experiments demonstrate that our model outperforms existing state-of-the-art methods in most cases, achieving higher predictive accuracy and better generalization. Our results validate the effectiveness of our approach in addressing the key challenges in time series forecasting, paving the way for more reliable and efficient predictive models in practical applications.
1 Introduction
@@
The field of time series forecasting has experienced significant advancements and found widespread applications in various real-world engineering tasks, such as traffic management [1], pandemic spread [2], and financial analysis [3]. From early models like GRU [4] and LSTM [5] to the more recent Transformer-based architectures [6, 7], a plethora of sophisticated models have emerged. For instance, TimesNet [8] and FreTS [9] approach data processing from a frequency domain perspective, while multi-scale models like Pyraformer [10] and Scaleformer [11] leverage intricate multi-scale and cross-dimensional dependency structures to significantly enhance performance on complex forecasting tasks.
@@
Howerver, existing models often require the introduction of additional parameters to capture multi-scale features, which increases computational complexity. Moreover, these models necessitate manual tuning of hyperparameters for different datasets, further complicating their application and limiting their generalization across diverse scenarios. More importantly, many models, including Autoformer [7] , struggle to effectively distinguish between signal and redundant features when dealing with data characterized by strong seasonality. Redundant features not only interferes with the learning process, leading to increased errors, but also obscures the true patterns within the data, making the predictions unreliable. This issue is particularly pronounced when traditional autocorrelation mechanisms break down due to redundant features interference, significantly diminishing the practical value of these models.
@@
To address these challenges, we proposes an innovative time series forecasting model designed to significantly improve predictive accuracy and broaden applicability. First, we employ wavelet transforms to preprocess the time series data using the Wavelet Decomposition Block (WDB) and the Wavelet Reconstruction Block (WRB). Wavelet transforms efficiently capture information from multiple scales, but this process may also introduce redundant features. To address this, we integrate the Information Filtering and Compression Block (IFCB), an information bottleneck module that compresses and extracts the most relevant features by filtering out redundant features from both detail and approximation coefficients. These components work together, significantly enhancing the model’s performance.
@@
  • This paper introduce WDB and WRB that efficiently capture multi-scale features in time series data without requiring manual hyperparameter tuning or the introduction of additional parameters.
@@
  • This paper utilize an information-theoretic approach by incorporating an IFCB to filter out redundant features from the data. This ensures that only the most predictive features are retained, enhancing the model’s accuracy.
@@
  • By integrating wavelet transforms for comprehensive feature extraction and the information bottleneck block for redundant features, our model achieves state-of-the-art performance, demonstrating superior predictive accuracy and generalization across diverse scenarios.
2 Related Work
2.1 Time Series Forecasting
@@
In recent years, the field of time series forecasting has advanced rapidly, giving rise to many efficient models. These models predominantly rely on multilayer perceptron (MLP) approaches like MSD-Mixer [12], TSMixer [13], DLinear [14]and FreTs [9], along with Transformer-based techniques such as PatchTST [15], Autoformer, and Informer [16], to perform predictions using sophisticated data transformation and learning mechanisms.
@@
Time series forecasting can be defined as follows: Given a historical sequence of observations $\ X=\{x_{1},x_{2},\ldots,x_{T}\}$, where $T$ is the lookback window length, the goal is to predict future values $\ Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}=F(X)$, where $H$ is the forecast window length and $F$ is the predictive model mapping the input sequence $X$ to the output sequence $Y$. However, time series data often exhibit complex, multi-scale characteristics. Conventional feature extraction methods, like standard pooling, require additional parameter adjustments or rely on manually designed pooling sizes, increasing model complexity and limiting flexibility. Wavelet transform, a classical signal processing technique [17], excels in multi-scale analysis. It decomposes signals into different frequency components, aiding in the understanding and prediction of intricate patterns.
@@
We propose a novel approach that combines wavelet transform with modern machine learning techniques for time series forecasting. The wavelet transform preprocesses the input data, decomposing it into detail coefficients and approximation coefficients at different scales. We believe this method effectively captures comprehensive features from the time series data.
2.2 Information Bottleneck Theory
@@
In recent years, with the growing intersection of information theory and deep learning research, many researchers have begun analyzing and optimizing neural networks from the perspective of information theory. In this trend, information bottleneck (IB) theory has gained particular attention [18, 19], finding wide application in emerging fields such as graph neural networks [20], multimodal learning [21], and recently in time series imputation [22]. IB theory helps improve model performance by optimizing information flow and processing [18].
@@
We believe that wavelet transformation can effectively decompose time series data into components that contain valuable information. However, these components also include a significant amount of redundant information. If not properly addressed, this redundant information can interfere with the model’s learning performance and predictive accuracy.
From the perspective of information theory, IB theory provides a mechanism to effectively retain critical information related to target variables while filtering out irrelevant or redundant information in the model’s intermediate representations. This approach enhances the model’s ability to capture key features. Furthermore, it strikes a delicate balance between compressing input information and maintaining predictive accuracy, ensuring that the model’s output remains precise and concise by:
$\max I(y;z)$ (1)
@@
3 TimeSieve
Figure 1: Overview of the TimeSieve Framework. The upper part illustrates the overall architecture of TimeSieve, which consists of WDB, WRB and IFCB. The input time series data, $T\times C$, is decomposed by WDB into coefficients $\pi_{a}$ and $\pi_{d}$, which are processed by IFCB. The WRB then reconstructs the data, followed by the predictor generating the corresponding forecast steps. The lower part provides a detailed visualization of the data flow within WDB and IFCB, with further visualization of IFCB available in Figure 2.
Figure 2: Overview of IFCB: This module consists of a reparameterization step and deep neural networks. Initially, the standard deviation and mean ($\mu_{z},\Sigma_{z}$) of the data are computed. The reparameterization is then performed using the equation $z=\mu_{z}+\Sigma_{z}\times\epsilon$. The input coefficients are processed through the first neural network to obtain $\mu_{z}$ and $\Sigma_{z}$. The reparameterized $z$ is then fed into the second neural network to produce the reconstructed coefficients $\hat{\pi_{i}}$.
In our study, we propose a novel time series forecasting model named TimeSieve. This model integrates the Information Filtering and Compression Block (IFCB) with wavelet transform method to enhance the accuracy of time series predictions.
We explain the whole framework first, and we explain WDB, WRB and IFCB in more detail after that.
@@ 3 TimeSieve
$[\pi_{a},\pi_{d}]=WDB(x(t))$ (2)
where $x(t)$ represents the input, the approximation coefficients $\pi_{a}$ represents the low-frequency trends, and the detail coefficients $\pi_{d}$ represents the high-frequency details, respectively.
However, the extracted multi-scale information may contain redundant features, which can adversely affect the model’s learning performance and predictive accuracy. To address this issue, we introduce the IFCB to filter the information. Additionally, we employ residual connections to ensure completely keeping the valuable information during the filtering process. The IFCB optimizes the information flow, retaining critical information while filtering out irrelevant or redundant information. The IFCB could be shown as following equation:
$\hat{\pi_{a}}=IFCB(\pi_{a}),\hat{\pi_{d}}=IFCB(\pi_{d})$ (3)
where $\hat{\pi_{a}}$, $\hat{\pi_{d}}$ represents the filtered coefficients $\pi_{a}$, $\pi_{d}$.
After filtering with the IFCB, we apply the Wavelet Reconstruction Block (WRB) to reconstruct the processed data back into the time domain. This step ensures that the features at different scales are fully utilized. Finally, this paper use an MLP as the prediction layer to make the final forecast on the processed time series data, specifically adjusting the output to match the desired forecast length. The prediction step is formulated as follows:
$Y=MLP(WRB(\hat{\pi_{i}})),i\in[a,d]$ (4)
@@
WDB and WRB. To effectively capture comprehensive feature information, we employ wavelet transformation to decompose time series data inspired by [23, 24], allowing us to analyze the data across multiple scales. Wavelet transformation allows us to break down the data into various frequency components, enabling us to analyze and extract features across multiple temporal scales. This multi-scale analysis is crucial for identifying patterns and trends that occur at different time resolutions.
WDB. For the input time series $x(t)\in\mathbb{R}^{T\times C}$, where $T$ represents the length of the lookback window and $C$ denotes the number of variates, the wavelet transform decomposes the series into approximation coefficients $\pi_{a}$ and detail coefficients $\pi_{d}$. The approximation coefficients $\pi_{a}$ capture the general trend or smooth component of the series, representing low-frequency information. In contrast, the detail coefficients $\pi_{d}$ capture the fine-grained, high-frequency information, highlighting local fluctuations and details within the series. This decomposition can be expressed as:
@@
$\phi(x)=\sum_{k=0}^{N-1}a_{k}\phi(2u-k),\psi(x)=\sum_{k=0}^{M-1}b_{k}\phi(2u-k)$ (6)
In the case of wavelet transform, $\phi(t)$ represents the low-pass filter, capturing the smooth, approximate part of the signal, and $\psi(t)$ represents the high-pass filter, capturing the detailed, high-frequency part of the signal. Here, $a_{k}$ and $b_{k}$ are the filter coefficients for the scaling and wavelet functions, respectively, $u$ is the variable over which the functions are defined, and $k$ is the index of the coefficients.
@@
IFCB. Motivated by the need to effectively filter out redundant features while retaining critical information, we introduce IFCB to achieve this goal, inspired by the information bottleneck principle [18, 19, 25]. In this section, we process the approximation coefficients $\pi_{a}$ and detail coefficients $\pi_{d}$ from the wavelet transform of the time series data, for convenience in subsequent explanations, we define $\pi_{i}$ and $i\in[a,d]$ to collectively represent $\pi_{a}$ and $\pi_{d}$. Additionally, We use $\hat{\pi_{i}}$ to represent the filtered form(s) of $\pi_{i}$. Initially, we define the input to the IFCB as $\pi_{i}\in\mathbb{R}^{T/2\times C}$ and the output to the IFCB as $\hat{\pi_{i}}\in\mathbb{R}^{T/2\times C}$.
In our model, both $\pi_{i}$ and $\hat{\pi_{i}}$ are treated as random variables, which given a joint distribution $p(i,\hat{i})$. The mutual information between these two random variables denoted as $I(\pi_{i};\hat{\pi_{i}})$, quantifies the amount of information shared by $pi_{i}$ and $\hat{\pi_{i}}$ Mutual information between joint distribution is defined as followed,
$I(\pi_{i};\hat{\pi_{i}})$ $=D_{kL}[p(i,\hat{i})\|p(i)p(\hat{i})]=\sum_{i\in\mathcal{I},\hat{i}\in\hat{\mathcal{I}}}{p\left(i,\hat{i}\right)\log\left(\frac{p\left(i,\hat{i}\right)}{p\left(i\right)p\left(\hat{i}\right)}\right)}$ (7)
$=\sum_{i\in\mathcal{I},\hat{i}\in\hat{\mathcal{I}}}{p\left(i,\hat{i}\right)\log\left(\frac{p\left(i|\hat{i}\right)}{p\left(i\right)}\right)}=H\left(\pi_{i}\right)-H\left(\pi_{i}|\hat{\pi_{i}}\right)$
where $D_{kL}[p\|q]$ is the Kullback-Liebler divergence of the distributions $p$ and $q$, and $H(P)$ and $H(P|Q)$ are the entropy and conditional entropy of $P$ and $Q$, repectively.
We define the intermediate hidden layer random variable $Z$, which follows the Markov chain principles, where $\pi_{i}\rightarrow Z\rightarrow\hat{\pi_{i}}$, the mutual information satisfies $I(\pi_{i};Z)\geq I(\pi_{i};\hat{\pi_{i}})$. Although there is an inevitable information loss in the transmission to the middle layer $Z$, notice that $Z$ retains the most critical information for predicting $\hat{\pi_{i}}$ while filtering out irrelevant features.
This property makes the definition of objective of IFCB reasonable, which is to maximize the mutual information between $Z$ and $\hat{\pi_{i}}$ while minimizing the information between $Z$ and $\pi_{i}$:
$\min(I(\pi_{i};Z)-\beta I(Z;\hat{\pi_{i}}))$ (8)
where $\beta$ is a trade-off parameter, which determines the weight of the minimal information constraint in the optimization process.
The optimization is carried out within a deep neural network, which is designed to minimize the information between $Z$ and $\pi_{i}$ by adjusting the network weights and biases through backpropagation and suitable regularization techniques.
Then optimization is performed within a deep neural network. Assuming that $p(z|i)$ follows a Gaussian distribution, and the following derivation is established:
$p(z|i)=\mathcal{N}(\mu(i;\theta_{\mu}),\Sigma(i;\theta_{\Sigma}))=\mathcal{N}(\mu_{z},\Sigma_{z})$ (9)
The parameter updating process posing a new challenge due to the stochastic property of the gradient calculations. To address the challenge, the reparameterization trick is introduced. The reparameterization is formulaed by:
$z=\mu_{z}+\Sigma_{z}\times\epsilon$ (10)
where $\epsilon$ is sampled from a standard Gaussian distribution.
Given that the study focuses on a regression task, we define the conditional probability $q(\hat{i}|z)$:
$q(\hat{i}|z)=e^{-||\hat{i}-D(z;\theta_{c})||}+C$ (11)
where $D$ is a decoder function parameterized by $\theta_{c}$, and $C$ is a constant. This formulation measures the fit of the model’s fitting by capturing the deviation of the predicted output $\hat{i}$ from the output of the decoder.
@@
where $\hat{\pi_{a}}$ and $\hat{\pi_{d}}$ are the filtered approximation and detail coefficients, respectively. Finally, the reconstructed time series is passed through a MLP for final prediction.
Algorithm 1 Pseudo code of TimeSieve
0:  Time series data $x(t)=\{x_{1},x_{2},\ldots,x_{T}\}$, true labels $Y=\{y_{T+1},y_{T+2},\ldots,y_{T+H}\}$
0:  Predicted output $\hat{Y}$
  Initialization: Initialize the model parameters, learning rate $\eta$, and total training epochs $E$
  Training:
  for $epoch\leftarrow 1$ to $E$ do
     $\pi_{a},\pi_{d}$leftarrow WDB($X(t)$)
     for $i\in\{a,d\}$ do
        $\hat{\pi_{i}},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(\pi_{i})$
     end for
     $X_{\text{reconstructed}}\leftarrow$ WRB($\hat{\pi_{a}},\hat{\pi_{d}}$)
     $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
     $\mathcal{L}_{o}\leftarrow\text{LossFunction}(\hat{Y}_{i},Y_{i})$
     for $i\in\{a,d\}$ do
        $\mathcal{L}_{IB_{I}}\leftarrow D_{KL}[\mathcal{N}(\mu_{zI},\Sigma_{zI})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z_{I})\,||\,p(z_{I}|i)]$
     end for
     $\mathcal{L}_{IB}\leftarrow\mathcal{L}_{IB_{A}}+\mathcal{L}_{IB_{D}}$
     $\mathcal{L}\leftarrow\mathcal{L}_{o}+\mathcal{L}_{IB}$
     $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
  end for
  Inference:
  $\pi_{a},\pi_{d}$leftarrow WDB($x(t)$)
  for $i\in\{a,d\}$ do
     $\hat{\pi_{i}},\mu_{z},\Sigma_{z}\leftarrow\text{IFCB}(\pi_{i})$
  end for
  $X_{\text{reconstructed}}\leftarrow$ WRB($\hat{\pi_{a}},\hat{\pi_{d}}$)
  $\hat{Y}_{i}\leftarrow\text{MLP}(X_{\text{reconstructed}})$
  return  final predictions $\hat{Y}$
@@
Loss Funtion. We now specify the model’s loss function. This function is crucial for training the model to minimize prediction errors and optimize the distribution parameters effectively. The loss is composed of the original loss component and the IFCB loss, as shown below:
$\mathcal{L}=\mathcal{L}_{o}+\mathcal{L}_{IB}=\mathcal{L}_{o}+D_{KL}[\mathcal{N}(\mu_{z},\Sigma_{z})\,||\,\mathcal{N}(0,I)]+D_{KL}[p(z)\,||\,p(z|i)]$ (13)